OpenStudy (anonymous):

What is the probability of getting tails each time if a coin is tossed 4 times?

OpenStudy (anonymous):

P(tails) = 0.5 P(tails 4 times in a row) = 0.5*0.5*0.5*0.5 = 1/16

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

If you look at the as the probability of getting 1 or more tail in 4 coin tosses, you would then calculate the probability of tossing 4 heads in a row and subracting that from 1. The probability fo tossing 4 heads is 1/2 * 1/2 * 1/2 *1/2 = 1/16. 1 - 1/16 = 15/16.

OpenStudy (anonymous):

oh ok thank u

OpenStudy (anonymous):

There are 3 literature books, 4 geography books, and 3 science books on a shelf. If 3 books are chosen at random one after the other, what is the probability that a literature book, a geography book, and a science book are selected if replacement does not take place?

OpenStudy (anonymous):

1/16